# Eigenvalue Gaps of the Laplacian of Random Graphs

### v1

@@ 2.2. No-gaps and no-structure delocalization
Our plan for Step 1 of the modified plan has two substeps
@@
The first type of result is called no-gaps delocalization, a notion pioneered by Rudelson and Vershynin in [RV16]. Note that in Random Matrix Theory, delocalization usually means that the eigenvectors are flat, not localized. For instance for the current centered Laplacian model it has been shown in [HL20] that all unit eigenvectors ${\mathbf{v}}$ (corresponding to the eigenvalues in the bulk) are completely delocalized in the sense that $\|{\mathbf{v}}\|_{\infty}\leq n^{{-1/2+o(1)}}$ with very high probability. So $\|{\mathbf{v}}_{I}\|_{2}$ is small over any set $I$ of size $o(n)$. The no-gaps delocalization, on the other hand, addresses the property that $\|{\mathbf{v}}_{I}\|_{2}$ cannot be small over any set $I$ of order $n$. We will cite below a very strong result of Rudelson-Vershynin in [RV16], that works for any general random non-symmetric, symmetric and skew-symmetric matrices $X_{n}$ (where $x_{ij}$ and $x_{ji}$ can depend on each other, but otherwise are independent, subgaussian of mean zero and variance one).
@@
There exist constants $C,C^{\prime},c$ such that the following holds. Let $\lambda_{n}\geq\frac{1}{n}$ and $s\geq C\lambda_{n}^{-7/6}n^{-1/6}+\exp(-c/\sqrt{\lambda_{n}})$. Then with probability at least $1-(C^{\prime}s)^{\lambda_{n}n}$, every unit eigenvector ${\mathbf{v}}$ of $X_{n}$
@@
While this result is extremely strong, the method of [RV16] does not seem to extend to (i) matrices where the diagonal entries depend on other entries (the way $\overline{L}_{n}$ does), and (ii) matrices of norm of order $\omega(\sqrt{n})$ – while here with high probability $\overline{L}_{n}$ has norm of order $\Theta(\sqrt{n\log n})$.
In this note we show the following analogue
@@
for some sufficiently large constant $A_{0}$, is bounded by $n^{-\omega(1)}$ as $n\to\infty$.
As a corollary, when $\lambda_{n}$ is of constant order, with high probability for any non-trivial unit eigenvector ${\mathbf{v}}$, not only is the total mass over $I$ not small, $\|{\mathbf{v}}_{I}\|_{2}\geq\frac{1}{(\log n)^{O(1)}}$, but it is also not small after shifting (affine delocalization)
@@
This translates into a statement about “incompressibility” of affine shifts of segments of eigenvectors. It is possible that the result continues to hold for $\inf_{a\in{\mathbb{R}}}\|{\mathbf{v}}_{I}-a1_{I}\|_{2}\geq\Theta_{\lambda_{n}}(1)$ and for $\lambda_{n}\ll\frac{1}{\log\log n}$ (with possibly much finer approximation than (5)) as in [RV16] but it seems $\frac{1}{(\log n)^{O(1)}}$ (for the mass contribution) and $\frac{1}{\log\log n}$ (for the lower bound of $\lambda_{n}$) are the limit of our proof. Theorem 2.3 applies to subsets $I$ of size larger than $n/\log\log n$ which is enough for our purposes, and also, the affine aspect is new and will play a key role for later parts of the paper. While the result is much weaker compared to Theorem 2.2, our matrix model, the random Laplacian, seems to be more involved and we will have to develop new ways to handle these difficulties. Furthermore, the behavior of eigenvectors for the Laplacian are fundamentally different than that of the non-symmetric or Wigner matrices. Even for a Laplacian matrix generated from gaussian random variables, the eigenvectors are not uniformly distributed over the sphere. In fact, simulations show that eigenvectors corresponding to large eigenvalues appear to be mildly localized.
Next, we will discuss Substep 1(ii), which is another innovative aspect of our note. To do this, we first introduce an important concept following Rudelson and Vershynin.
@@
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}})=\inf\left\{\theta>0:{\operatorname{dist}}(\theta{\mathbf{x}},{\mathbb{Z}}^{n})<\min\{\gamma\|\theta{\mathbf{x}}\|_{2},\kappa\}\right\}$
@@
One of the highlights of the paper is to show that not only are the eigenvectors no-gaps delocalized, but also “no-structure delocalized”.
@@
Let $A>0$ be given. Let $n_{0}=\lambda_{n}n$ where
@@
Here the bound $\frac{1}{\sqrt{\log\log n}}\leq\lambda_{n}$ is likely to be improvable to $\frac{1}{\log\log n}\leq\lambda_{n}$ as in Theorem 2.3 but we will not focus on this.
2.4. Approximated eigenvetors and more general results
@@
We call the above ${\mathbf{v}}$ approximated non-trivial eigenvectors.
@@
Theorems 2.3 and 2.4 hold for approximated eigenvectors from (7) of $\overline{L}^{\prime}_{n-1}$ (or of $\overline{L}_{n-1}$).
@@ 2.5. Overcrowding estimate
One of the main obstacles to rule out the event of having $\|{\mathbf{v}}_{I}\|_{2}=o(1)$ or $\mathbf{LCD}({\mathbf{v}}_{I})$ small is that we have no information on the remaining vector ${\mathbf{v}}_{I^{c}}$. For the sake of discussion, let us assume that $X_{n}{\mathbf{v}}=0$. We can rewrite this as $M_{11}{\mathbf{v}}_{I}+M_{21}{\mathbf{v}}_{I^{c}}=0$ and $M_{12}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}=0$, where
@@
Given ${\mathbf{v}}_{I}$ from some finite set for which we can control the size, conditioning on $M_{21},M_{12}$ and $M_{22}$ we can solve for ${\mathbf{v}}_{I^{c}}$ from $M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}=0$, that ${\mathbf{v}}_{I^{c}}=-M_{22}^{{-1}}M_{21}{\mathbf{v}}_{I}$. Substitute this into the first equation, we hence obtain
@@
One of the key problems of this approach, taking into account that we are working with asymptotic approximations rather than with identities, is that we have to control the norm of $\|M_{22}^{-1}\|_{2}$. This seems to be a very hard problem, where it is only known recently from [NW24] that $M_{22}$ is invertible with high probability. To avoid this difficulty, we will not work with $M_{22}$ directly, but a near-square rectangular submatrix $M_{22}^{\prime}$ of it, where we can control the non-zero least singular values quite effectively. One drawback of this modification is that we can no longer solve for ${\mathbf{v}}_{I^{c}}$, but for only (and approximately) a major part of it. However, this problem can be resolved by again passing to another approximated eigenvector. We now state the main result on the strong invertibility of $M_{22}^{\prime}$ in the form of $L_{n}$.
@@
Let $L_{n}$ be the Laplacian corresponding an Erdős-Rényi graph $G(n,p)$. There exist constants $C\geq 1$, $c>0$, depending only on $p$ such that
@@
This result, which can be seen as an analogue of Theorem 1.7 in [Ngu18] where iid and symmetric matrices were considered, is of its own interest. Interestingly, as we can see, the result works for all choices of deterministic matrices $F$, and hence works for $\overline{L}_{n}$ as well  [^4] The proof automatically extends to the model $\overline{L}^{\prime}_{n}$..
The dependence on $k$ in the exponential probability bound in Theorem 2.6 can likely be improved to $k^{2}$ as in [Ngu18]. However, we do not pursue this extension further in this paper as for our application, it suffices that when $k=\Theta(n/\log^{C}n)$, the probability bound is super exponentially small.
@@ 2.6. Small coordinates of eigenvectors
As first observed in [NTV17], a direct consequence of controlling the structure of the eigenvectors of a principle minor is that the eignevectors of the Laplacian cannot have many coordinates of small size (in fact, we show they cannot have more than one zero coordinate). The zero-coordinates of Laplacian eigenvectors play a special role in nodal domains in spectral graph theory and the physics of mechanical systems, where they are referred to as soft nodes [DLL11, BykgHL+04, CKK19, CK23]. For dynamical systems on graphs governed by the Laplacian, soft nodes are of interest theoretically and practically because they are immune to forcing and damping at that node [CKS12].
@@
Naturally, we expect that the reality is that there are no small coordinates with high probability, as was shown for Wigner matrices in [NTV17].
@@ 2.7. Further Remarks
Additionally, it seems to be an important problem to extend these results to Laplacians of sparse graphs $G(n,p)$, where $p\rightarrow 0$ as $n\rightarrow\infty$, as well as to random $d$-regular graphs (for either fixed $d$ or $d\to\infty$ with $n$). Finally, it is of interest to study the law of the minimum gap of these random Laplacians, with a focus on demonstrating universality.
@@ 3. Supporting lemmas
Lemma 3.1 (Theorem 6 of [NY17]).
Let $A:{\mathbb{R}}^{n}\to{\mathbb{R}}^{k}$ of full rank $k$. Then for every $11\leq l\leq k-$ 1, there exists a $I\subseteq\{1,\ldots,n\}$ of size $l$ such that if $A_{I}$ denotes $A\mid_{{\mathbb{R}}^{I}}$, then
$\sigma_{l}(A_{I})\geq c\max_{r\in\{l+1,\ldots,k\}}\sqrt{\frac{(r-l)\sum_{i=r}^{n}\sigma_{i}(A)^{2}}{nr}}$
@@
Next we need the following elementary fact for the centered Laplacian matrices.
@@
In particular, with probability at least $1-\exp(-\Theta(\log^{2}n))$, all submatrices of $\overline{L}_{n}$ have norm, say, bounded by $\sqrt{n}\log n$.
Proof.
(of Lemma 3.2) By the triangle inequality
@@
We can then bound $\|D_{n}-{\mathbf{E}}D_{n}\|_{2}$ and $\|A_{n}-{\mathbf{E}}A_{n}\|_{2}$ separately. Alternatively, it follows from [Ban18, Theorem 3.2] that the median of $\|\overline{L}_{n}\|_{2}$ has order $\Theta(\sqrt{n\log n})$. Then by Talagrand concentration theorem, as $\|\overline{L}_{n}\|_{2}$ is convex and $\sqrt{n}$-Lipschitz with respect to the .$\|.\|_{HS}$-norm, we have
@@
Occasionally, in many proofs we will need the following tensorization Lemma to transform anti-concentration bound from independent random variables to random vectors ([RV08, Lemma 3.4]).
@@
Proof.
(of Lemma 3.3) Let $\delta\geq\delta_{0}$, we have the following by Chebychev’s inequality:
@@
Integrating the tail bound we have
@@
  • For any set $S$ of size $\lfloor c_{0}n\rfloor$, there exists one $i\in S$ such that $|x_{i}|\geq(c_{1}/c_{0})^{{1/2}}/\sqrt{n}$. It thus follows that for all but $\lfloor c_{0}n\rfloor-1$ indices $i$ we have $|x_{i}|\geq(c_{1}/c_{0})^{{1/2}}/\sqrt{n}$.
@@
The next Lemma connects the notion of compressible vectors and $\mathbf{LCD}$ together.
@@
It is worth noting that the smaller $\gamma$ is, the larger the $\mathbf{LCD}_{\kappa,\gamma}$ becomes.
Proof.
(of Lemma 3.5) For each $k\in\mathrm{spread}({\mathbf{x}})$,
@@
$\mathbf{LCD}_{\kappa,\gamma\|{\mathbf{x}}\|_{2}/\|{\mathbf{x}}^{\prime}\|_{2}}(\frac{{\mathbf{x}}^{\prime}}{\|{\mathbf{x}}^{\prime}\|_{2}})\leq\frac{\|{\mathbf{x}}^{\prime}\|_{2}}{\|{\mathbf{x}}\|_{2}}\mathbf{LCD}_{\kappa,\gamma}(\frac{{\mathbf{x}}}{\|{\mathbf{x}}\|_{2}}).$
Proof.
(of Fact 3.6) It suffices to assume $\|{\mathbf{x}}\|_{2}=1$. Note that for any $D>0$, if ${\operatorname{dist}}(D{\mathbf{x}},{\mathbb{Z}}^{n})\leq\min(\gamma D,\kappa)$ then trivially
@@
Theorem 3.7 (Small ball probability via LCD, [RV08]).
Let $\xi$ be a sub-gaussian random variable of mean zero and variance one, and let $\xi_{1},\cdots,\xi_{n}$ be iid copies of $\xi$. Consider a vector ${\mathbf{x}}\in{\mathbb{R}}^{n}$ which satisfies $\left\|{\mathbf{x}}\right\|\geq 1$. Then, for every $\kappa>0$ and $\gamma\in(0,1)$, and for
@@
Finally, we need the following important results, whose proof will be given for the reader’s convenience.
@@
  • (Nets of vectors having $\mathbf{LCD}$ belonging to $(D_{0},2D_{0}]$, [RV09, Lemma 4.7][NTV17, Lemma 11.3]) Assume that $\kappa/D_{0}\leq 1/2$. Let $S_{\alpha,D_{0}}$ be the set of ${\mathbf{x}}\in{\mathbb{R}}^{m}$ with $\|{\mathbf{x}}\|=\alpha$ and that
@@
    Then there exists a $(2\alpha\kappa/D_{0})$-net of $S_{D_{0}}$ of cardinality at most $(C_{0}D_{0}/\sqrt{m})^{m}$, where $C_{0}$ is an absolute constant. Consequently, there exists a $(2\kappa/D_{0})$-net of $S_{\alpha,D_{0}}$ of cardinality at most $(C_{0}(\alpha+1)D_{0}/\sqrt{m})^{m}$.
  • (Nets of unit vectors having $\mathbf{LCD}$ smaller than $D$, [NTV17, Lemma 11.4]) Assume that $2\kappa\leq D_{0}\leq D$. Then the set $S_{1,D_{0}}$ has a $(2\kappa/D)$-net of cardinality at most $(C_{0}D/\sqrt{m})^{m}$ for some absolute constant $C_{0}$.
@@
Proof.
(of Lemma 3.8) Let us first focus on the net of $S_{\alpha,D_{0}}$. For ${\mathbf{x}}\in S_{\alpha,D_{0}}$ and $\|{\mathbf{x}}\|_{2}=\alpha$, denote
@@
This implies that $\|p\|\approx D({\mathbf{x}})$, more precisely
@@
For the consequence, it suffices to assume $\alpha\geq 1$. As we can cover $S_{\alpha,D_{0}}$ by $(C_{0}D_{0}/\sqrt{m})^{m}$ balls of radius $2\alpha\kappa/D_{0}$, we can then cover these balls by smaller balls of radius $2\kappa/D$, the number of such small balls is at most $(O(\alpha))^{m}$. Thus there are at most $(C_{0}\frac{\alpha D_{0}}{\sqrt{m}})^{m}$ balls of radius $2\kappa/D$ in total.
We can justify the second statement similarly. By the first part, one can cover $S_{1,D_{0}}$ by $(C_{0}D_{0}/\sqrt{m})^{m}$ balls of radius $2\kappa/D_{0}$. We then cover these balls by smaller balls of radius $2\kappa/D$, the number of such small balls is at most $(O(D/D_{0}))^{m}$. Thus there are at most $(O(D/\sqrt{m}))^{m}$ balls in total. ∎
@@
We first introduce an important lemma (whose proof is delayed to the end of the section) to control the distance of a random vector with non i.i.d. entries to a subspace.
@@
Let ${\mathbf{v}}\in{\mathbb{R}}^{n}$ be a random vector whose entries are independent (not necessarily identically distributed) and ${\mathbf{v}}_{i}$ have mean zero, ${\operatorname{Var}}({\mathbf{v}}_{i})\geq 1$ and $|{\mathbf{v}}_{i}|\leq T$ for some parameter $T$ with probability one. Then if $P_{H^{\perp}}$ is a deterministic orthogonal projection in ${\mathbb{R}}^{n}$ onto a subspace of dimension $k$, there exist constants $C,c,c^{\prime}$ depending on $p$ such that for any $0<t<1/2,$
@@
Proof.
(of Theorem 2.6) We follow the proof strategy in [Ngu18]. Observe that as $L_{n}+F$ is symmetric, $\sigma_{n-k+1}(L_{n}+F)\leq\frac{ck}{\sqrt{n}}$ is equivalent to the event that there exists an $i$ such that

### v2

@@ 2.2. No-Gaps and No-Structure Delocalization
Step 1 of our modified plan has two substeps.
@@
The first type of result is called no-gaps delocalization, a notion pioneered by Rudelson and Vershynin in [60]. Note that in random matrix theory, delocalization usually means that the eigenvectors are flat, not localized. For instance, for the current centered Laplacian model, it has been shown in [39] that all unit eigenvectors ${\mathbf{v}}$ corresponding to the eigenvalues in the bulk are completely delocalized in the sense that $\|{\mathbf{v}}\|_{\infty}\leq n^{{-1/2+o(1)}}$ with very high probability. So $\|{\mathbf{v}}_{I}\|_{2}$ is small over any set $I$ of size $o(n)$. The no-gaps delocalization, on the other hand, addresses the property that $\|{\mathbf{v}}_{I}\|_{2}$ cannot be small over any set $I$ of order $n$. We will cite below a very strong result of Rudelson-Vershynin in [60], that works for any general random non-symmetric, symmetric and skew-symmetric matrices $X_{n}$ (where $x_{ij}$ and $x_{ji}$ can depend on each other, but otherwise are independent, subgaussian of mean zero and variance one).
@@
There exist constants $C,C^{\prime},c$ such that the following holds. Let
$\lambda_{n}\geq\frac{1}{n}\text{ and }s\geq C\lambda_{n}^{-7/6}n^{-1/6}+\exp(-c/\sqrt{\lambda_{n}}).$
Then with probability at least $1-(C^{\prime}s)^{\lambda_{n}n}$, every unit eigenvector ${\mathbf{v}}$ of $X_{n}$
@@
Although this result is extremely strong, the method of [60] does not seem to extend to matrices where the diagonal entries depend on other entries (the way $\overline{L}_{n}$ does), nor to matrices of norm of order $\omega(\sqrt{n})$ — in our setting, $\overline{L}_{n}$ has norm of order $\Theta(\sqrt{n\log n})$ with high probability.
In this paper, we show the following analogue
@@
For some sufficiently large constant $A_{0}$, the probability with respect to $\overline{L}^{\prime}_{n-1}$ (or $\overline{L}_{n-1}$) that it has a unit eigenvector ${\mathbf{v}}$ and an index set $I\subset[n]$ of size $n_{0}$ such that ${\mathbf{v}}\perp\mathbf{1}$ and
@@
is bounded by $n^{-\omega(1)}$ as $n\to\infty$.
As a corollary, when $\lambda_{n}$ is of constant order, with high probability, for any non-trivial unit eigenvector ${\mathbf{v}}$, not only is the total mass over $I$ not small, $\|{\mathbf{v}}_{I}\|_{2}\geq\frac{1}{(\log n)^{O(1)}}$, but it is also not small after shifting (affine delocalization):
@@
This translates into a statement about the “incompressibility” of affine shifts of segments of eigenvectors. It is possible that the result continues to hold for $\inf_{a\in{\mathbb{R}}}\|{\mathbf{v}}_{I}-a1_{I}\|_{2}\geq\Theta_{\lambda_{n}}(1)$ and for $\lambda_{n}\ll\frac{1}{\log\log n}$ (with possibly much finer approximation than (5)) as in [60], but it seems that $\frac{1}{(\log n)^{O(1)}}$ (for the mass contribution) and $\frac{1}{\log\log n}$ (for the lower bound of $\lambda_{n}$) are the limits of our proof. Theorem 2.3 applies to subsets $I$ of size larger than $n/\log\log n$ which is enough for our purposes, and also, the affine aspect is new and will play a key role for later parts of the paper. While the result is much weaker compared to Theorem 2.2, our matrix model, the random Laplacian, is more involved and we develop new ways to handle these difficulties. Furthermore, the behavior of eigenvectors of the Laplacian is fundamentally different than that of the non-symmetric or Wigner matrices. Even for a Laplacian matrix generated from Gaussian random variables, the eigenvectors are not uniformly distributed over the sphere. In fact, simulations show that eigenvectors corresponding to large eigenvalues appear to be mildly localized.
Next, we will discuss Substep 1(ii), which is another innovative aspect of our paper. To do this, we first introduce an important concept following Rudelson and Vershynin.
@@
$\mathbf{LCD}_{\kappa,\gamma}({\mathbf{x}})=\inf\left\{\theta>0:{\operatorname{dist}}(\theta{\mathbf{x}},{\mathbb{Z}}^{n})<\min\{\gamma\|\theta{\mathbf{x}}\|_{2},\kappa\}\right\}.$
For notational convenience, whenever we refer to $\mathbf{LCD}_{\kappa,\gamma}({{\mathbf{x}}})$ for any non-zero vector ${\mathbf{x}}\in{\mathbb{R}}^{n}$, we always mean $\mathbf{LCD}_{\kappa,\gamma}\left(\frac{{\mathbf{x}}}{\|{\mathbf{x}}\|_{2}}\right)$.
@@
One of the highlights of the paper is to show that, not only are the eigenvectors of the Laplacian no-gaps delocalized, but also “no-structure delocalized”.
@@
Let $A>0$ be given. Let $n_{0}=\lambda_{n}n$, where
@@
Here the bound $\frac{1}{\sqrt{\log\log n}}\leq\lambda_{n}$ can likely be improved to $\frac{1}{\log\log n}\leq\lambda_{n}$ as in Theorem 2.3, but we will not dwell on this point here.
2.4. Approximate Eigenvectors and More General Results
@@
We call the above ${\mathbf{v}}$ an approximate non-trivial eigenvector.
@@
Theorems 2.3 and 2.4 hold for the approximate eigenvectors from (7) of $\overline{L}^{\prime}_{n-1}$ (or of $\overline{L}_{n-1}$).
@@ 2.5. Overcrowding estimate
One of the main obstacles in ruling out the event that $\|{\mathbf{v}}_{I}\|_{2}=o(1)$ or the event that $\mathbf{LCD}({\mathbf{v}}_{I})$ is small is that we have no information on the remaining vector ${\mathbf{v}}_{I^{c}}$. For the sake of discussion, let us assume that $X_{n}{\mathbf{v}}=0$. We can rewrite this as $M_{11}{\mathbf{v}}_{I}+M_{21}{\mathbf{v}}_{I^{c}}=0$ and $M_{12}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}=0$, where
@@
Given ${\mathbf{v}}_{I}$ from some finite set, for which we can control the size, conditioning on $M_{21},M_{12}$ and $M_{22}$ we can solve for ${\mathbf{v}}_{I^{c}}$ from $M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}=0$, that ${\mathbf{v}}_{I^{c}}=-M_{22}^{{-1}}M_{21}{\mathbf{v}}_{I}$. Substituting this into the first equation, we hence obtain
@@
One of the key problems of this approach, taking into account that we are working with asymptotic approximations rather than with identities, is that we have to control the norm of $\|M_{22}^{-1}\|_{2}$. This seems to be a very hard problem, where it is only known recently from [55] that $M_{22}$ is invertible with high probability. To avoid this difficulty, we will not work with $M_{22}$ directly, but a near-square rectangular submatrix $M_{22}^{\prime}$ of it, where we can control the non-zero least singular values quite effectively. One drawback of this modification is that we can no longer solve for ${\mathbf{v}}_{I^{c}}$, but for only (and approximately) a major part of it. However, this problem can be resolved by again passing to another approximate eigenvector. We now state the main result on the strong invertibility of $M_{22}^{\prime}$ in the form of $L_{n}$.
@@
Let $L_{n}$ be the Laplacian corresponding to an Erdős-Rényi graph $G(n,p)$. There exist constants $C\geq 1$, $c>0$, depending only on $p$ such that
@@
This result, which can be seen as an analogue of Theorem 1.7 in [53] where i.i.d. and symmetric matrices were considered, is interesting in its own right. Interestingly, as we can see, the result works for all choices of deterministic matrices $F$, and hence works for $\overline{L}_{n}$ as well  [^4] The proof automatically extends to the model $\overline{L}^{\prime}_{n}$..
The dependence on $k$ in the exponential probability bound in Theorem 2.6 can likely be improved to $k^{2}$ as in [53]. However, we do not pursue this extension further in this paper since, for our application, it suffices that when $k=\Theta(n/\log^{C}n)$, the probability bound is super exponentially small.
@@ 2.6. Small coordinates of eigenvectors
As first observed in [52], a direct consequence of controlling the structure of the eigenvectors of a principal minor is that the eigenvectors of the Laplacian cannot have many coordinates of small size (in fact, we show they cannot have more than one zero coordinate). The zero coordinates of Laplacian eigenvectors play a special role in nodal domains in spectral graph theory and the physics of mechanical systems, where they are referred to as soft nodes [25, 10, 14, 15]. For dynamical systems on graphs governed by the Laplacian, soft nodes are of interest theoretically and practically because they are immune to forcing and damping [16].
@@
Naturally, we expect that there are no small coordinates with high probability, as was shown for Wigner matrices in [52].
@@ 2.7. Further Remarks
Additionally, it seems to be an important problem to extend these results to Laplacians of sparse graphs $G(n,p)$, where $p\rightarrow 0$ as $n\rightarrow\infty$, as well as to random $d$-regular graphs (for either fixed $d$ or $d\to\infty$ with $n$). Finally, a natural direction for future research is to study the law of the minimum gap of these random Laplacians, with a focus on demonstrating universality.
@@ 3. Supporting Lemmas
Lemma 3.1 (Theorem 6 of [51]).
Let $A:{\mathbb{R}}^{n}\to{\mathbb{R}}^{k}$ of full rank $k$. Then for every $11\leq l\leq k-$ 1, there exists an $I\subseteq\{1,\ldots,n\}$ of size $l$ such that if $A_{I}$ denotes $A\mid_{{\mathbb{R}}^{I}}$, then
$\sigma_{l}(A_{I})\geq c\max_{r\in\{l+1,\ldots,k\}}\sqrt{\frac{(r-l)\sum_{i=r}^{n}\sigma_{i}(A)^{2}}{nr}},$
@@
Next, we need the following elementary fact for the centered Laplacian matrices.
@@
In particular, with probability at least $1-\exp(-\Theta(\log^{2}n))$, all submatrices of $\overline{L}_{n}$ have norm bounded by $\sqrt{n}\log n$.
Proof of Lemma 3.2.
By the triangle inequality
@@
We can then bound $\|D_{n}-{\mathbf{E}}D_{n}\|_{2}$ and $\|A_{n}-{\mathbf{E}}A_{n}\|_{2}$ separately. Alternatively, it follows from [6, Theorem 3.2] that the median of $\|\overline{L}_{n}\|_{2}$ has order $\Theta(\sqrt{n\log n})$. Then by Talagrand concentration theorem, as $\|\overline{L}_{n}\|_{2}$ is convex and $\sqrt{n}$-Lipschitz with respect to the .$\|.\|_{HS}$-norm, we have
@@
Occasionally, in many proofs we will need the following tensorization lemma to transform anti-concentration bounds from independent random variables to random vectors ([58, Lemma 3.4]).
@@
Proof of Lemma 3.3.
Let $\delta\geq\delta_{0}$, we have the following by Chebyshev’s inequality:
@@
Integrating the tail bound, we have
@@
  • For any set $S$ of size $\lfloor c_{0}n\rfloor$, there exists an $i\in S$ such that $|x_{i}|\geq(c_{1}/c_{0})^{{1/2}}/\sqrt{n}$. It thus follows that for all but $\lfloor c_{0}n\rfloor-1$ indices $i$ we have $|x_{i}|\geq(c_{1}/c_{0})^{{1/2}}/\sqrt{n}$.
@@
The next lemma connects the notions of compressible vectors and $\mathbf{LCD}$.
@@
It is worth noting that as $\gamma$ decreases, $\mathbf{LCD}_{\kappa,\gamma}$ increases.
Proof of Lemma 3.5.
For each $k\in\mathrm{spread}({\mathbf{x}})$,
@@
$\mathbf{LCD}_{\kappa,\gamma\|{\mathbf{x}}\|_{2}/\|{\mathbf{x}}^{\prime}\|_{2}}\left(\frac{{\mathbf{x}}^{\prime}}{\|{\mathbf{x}}^{\prime}\|_{2}}\right)\leq\frac{\|{\mathbf{x}}^{\prime}\|_{2}}{\|{\mathbf{x}}\|_{2}}\mathbf{LCD}_{\kappa,\gamma}\left(\frac{{\mathbf{x}}}{\|{\mathbf{x}}\|_{2}}\right).$
Proof of Fact 3.6.
It suffices to assume $\|{\mathbf{x}}\|_{2}=1$. Note that for any $D>0$, if ${\operatorname{dist}}(D{\mathbf{x}},{\mathbb{Z}}^{n})\leq\min(\gamma D,\kappa)$ then trivially
@@
Theorem 3.7 (Small ball probability via LCD, [58]).
Let $\xi$ be a subgaussian random variable of mean zero and variance one, and let $\xi_{1},\cdots,\xi_{n}$ be i.i.d. copies of $\xi$. Consider a vector ${\mathbf{x}}\in{\mathbb{R}}^{n}$ which satisfies $\left\|{\mathbf{x}}\right\|\geq 1$. Then, for every $\kappa>0$ and $\gamma\in(0,1)$, and for
@@
Finally, we need the following important results, the proofs of which will be given for the reader’s convenience.
@@
  • (Nets of vectors having $\mathbf{LCD}$ belonging to $(D_{0},2D_{0}]$, [59, Lemma 4.7][52, Lemma 11.3]) Assume that $\kappa/D_{0}\leq 1/2$. Let $S_{\alpha,D_{0}}$ be the set of ${\mathbf{x}}\in{\mathbb{R}}^{m}$ with $\|{\mathbf{x}}\|=\alpha$ and that
@@
    Then there exists a $(2\alpha\kappa/D_{0})$-net of $S_{\alpha,D_{0}}$ of cardinality at most $(C_{0}D_{0}/\sqrt{m})^{m}$, where $C_{0}$ is an absolute constant. Consequently, there exists a $(2\kappa/D_{0})$-net of $S_{\alpha,D_{0}}$ of cardinality at most $(C_{0}(\alpha+1)D_{0}/\sqrt{m})^{m}$.
  • (Nets of unit vectors having $\mathbf{LCD}$ smaller than $D$, [52, Lemma 11.4]) Assume that $2\kappa\leq D_{0}\leq D$. Then the set $S_{1,D_{0}}$ has a $(2\kappa/D)$-net of cardinality at most $(C_{0}D/\sqrt{m})^{m}$ for some absolute constant $C_{0}$.
@@
Proof of Lemma 3.8.
Let us first focus on the net of $S_{\alpha,D_{0}}$. For ${\mathbf{x}}\in S_{\alpha,D_{0}}$ and $\|{\mathbf{x}}\|_{2}=\alpha$, denote
@@
This implies that $\|p\|\approx D({\mathbf{x}})$. More precisely, it implies
@@
For the second part of the first statement, it suffices to assume that $\alpha\geq 1$. As we can cover $S_{\alpha,D_{0}}$ by $(C_{0}D_{0}/\sqrt{m})^{m}$ balls of radius $2\alpha\kappa/D_{0}$, we can then cover these balls by smaller balls of radius $2\kappa/D$, the number of such small balls is at most $(O(\alpha))^{m}$. Thus there are at most $(C_{0}\frac{\alpha D_{0}}{\sqrt{m}})^{m}$ balls of radius $2\kappa/D$ in total.
We can justify the second statement similarly. By the first part, one can cover $S_{1,D_{0}}$ by $(C_{0}D_{0}/\sqrt{m})^{m}$ balls of radius $2\kappa/D_{0}$. We then cover these balls by smaller balls of radius $2\kappa/D$; the number of such small balls is at most $(O(D/D_{0}))^{m}$. Thus, there are at most $(O(D/\sqrt{m}))^{m}$ balls in total. ∎
@@
We first introduce an important lemma (whose proof is delayed to the end of the section) to control the distance of a random vector with non-i.i.d. entries to a subspace.
@@
Let ${\mathbf{v}}\in{\mathbb{R}}^{n}$ be a random vector whose entries are independent (not necessarily identically distributed) and ${\mathbf{v}}_{i}$ have mean zero, ${\operatorname{Var}}({\mathbf{v}}_{i})\geq 1$ and $|{\mathbf{v}}_{i}|\leq T$ for some parameter $T$ with probability one. Then, if $P_{H^{\perp}}$ is a deterministic orthogonal projection in ${\mathbb{R}}^{n}$ onto a subspace of dimension $k$, there exist constants $C,c,c^{\prime}$ depending on $p$ such that for any $0<t<1/2,$
@@
Proof of Theorem 2.6.
We follow the proof strategy in [53]. Observe that as $L_{n}+F$ is symmetric, $\sigma_{n-k+1}(L_{n}+F)\leq\frac{ck}{\sqrt{n}}$ is equivalent to the event that there exists an $i$ such that